OpenStudy (anonymous):

if \[f:\mathbb{R} \rightarrow \mathbb{R}\] with \[ f(x) :=(sin(x))^{2}\].Then its \[\forall x \in \mathbb{R}: f(x) = f(x+\pi)\]. TRUE OR FALSE ?

OpenStudy (anonymous):

sin(x+Pi) =-sin(x)

OpenStudy (anonymous):

and (-y)^2 = y^2

OpenStudy (anonymous):

true cos when you square (-sin x)^2=(sinx)^2

OpenStudy (fifciol):

what does f(x) look like? It's something like that: What is the period? Well, for sinx it is 2pi but if you have sin^2x all the negative values become positive values and the period becomes shorter by an amount of 2 so the statement is true